# Resonant phonons: Localization in a structurally ordered crystal

### v1

@@ I Introduction
Distinct from Bragg scattering, a recently discovered mechanism that has been shown to cause transformative changes to the phonon band structure is engineered intrinsic atomic-scale local resonances [39]  [^1] The notion of intrinsic local resonances has been introduced in the context of electromagnetics [94, 95] and acoustics [72, 96, 97]. Unlike thermal phonons, these applications are practically described as linear wave problems where the physical behaviour of interest is dominated by excitations at a single frequency or a narrowband frequency range.. The presence of sub-MFP vibrating substructures generates a multitude of wave-vector-independent resonant phonons (also described as vibrons [41]) that couple and hybridize with the underlying traveling phonons in a host medium [42]. Nanostructures exhibiting this behavior are referred to as nanophononic metamaterials (NPMs); see, for example, Refs. [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. Unlike the Bragg scatterers in an NPC, the nanoresonators in an NPM are not necessarily periodically spaced, which enhances the robustness of the mechanism. Thermal conductivity reduction by nanoscale local resonances has recently been experimentally demonstrated in suspended silicon membranes with a random array of GaN nanopillars standing on the surface [57]. This nanopillars-on-a-membrane configuration also yielded a decoupling between the thermal and electrical properties, which is highly advantageous for thermolectrics [58, 59].
@@
Here we reveal--for the first time--the appearance of spatial phonon localization along the direction of transport in this class of nanostructures. We provide evidence of this phenomenon by direct explicit observations from room-temperature atomic-scale simulations of silicon membranes with resonating silicon nanopillars standing on the surface. These simulations involve frequency-targeted wave-packet excitations that are applied over a thermalized atomic environment at equilibrium, where the full scope of phonon-phonon scattering events is accounted for  [^2] This contrasts sharply with the localization of linear elastic waves due to local resonances at the macroscopic scale [98].. The form of phonon localization we highlight occurs in the base membrane across multiple unit cells, which is distinctly different from the observation of a large concentration of phonon motion in a nanopillar in a single unit cell as described by mode shapes obtained by harmonic eigenvalue analysis--with the latter often also referred to as phonon localization in the literature [43, 47, 42]. The phenomenon we report, which we refer to as resonant localization, has profound impact on the thermal behavior and other condensed matter properties in general, and its assessment along the direction of transport enables direct comparison with AL and with coherent phenomena in other types of ordered nanostructures. (1) Unlike other coherent phonon mechanisms, this distinct phenomenon manifests across all the phonon spectrum, including low frequencies smaller than the mobility edge of a randomly perturbed lattice (AL) [61, 62, 16] or the Bragg limit corresponding to the periodicity of ordered nanostructure features (SL and NPCs) [42]. (2) Furthermore, unlike AL it is fully deterministically predictable in both space and time. (3) Another unique aspect of the resonant localization mechanism is that it is associated with minimal energy exchange with external sources at the resonant frequency bands.
@@ II Phonon Band Structures
Moving to NPMs, Fig. 1d shows the phonon band structure for a resonant unit cell in the form of a membrane with a central squared nanopillar having the same height as the membrane. As seen in previous studies [44, 42], both the LD and MD-SED calculations show in the dispersion diagram the hybridization of numerous phonon modes in this geometry, distinguishable as horizontal lines associated with resonances coupled with the underlying host medium’s wave-vector-dependent propagating modes. The mode shapes of eigenstates populating these horizontal lines (and parts of the associated avoided crossing regions) indicate a concentration of the atomic kinetic energy in the nanoresonator portion of the unit cell, e.g., inside the nanopillar portion of a nanopillared membrane [43, 47, 42, 64]. In contrast to the loss of band gaps in NPCs, band shifts stemming from anharmonicity do not disrupt the presence of the local resonances and their ability to couple with propagating waves, even at room temperature--as shown for the silicon NPM of Fig. 1d. The occurrence of the local resonances may in principle be precisely synthesized to take place across the full phonon spectrum. Indeed, it is also observed in Fig. 1d that the resonance hybridizations take place across the full frequency range displayed, including at low frequencies below the Bragg limit associated with the size of the NPM unit cell. In principle, the number and specific frequencies of these resonances are a characteristic of the nanopillar geometry and choice of material, and are therefore controllable by design. The number of local resonances is three times the number of atoms in a unit nanoresonator, and bandwidth of each resonant mode is dependent on its dissipation properties [47].
@@ III Deterministic localization via local resonances
Figure 4: Wave-packet envelope evolution at room temperature. Maximum phonon kinetic energy measured around the excitation region due to wave-packet propagation at the resonant frequency (b) or non-resonant frequencies (a,c) in NPM. The kinetic energy is represented in logarithmic scale, and the dashed lines indicate exponential fits to the envelope relaxation. This fit is unsuccessful under resonant conditions (b), thus indicating unusual phonon propagation and relaxation beyond the particle-like scattering picture provided by the relaxation time approximation.
@@
To further explore the resonant localization phenomenon, it is illustrative to analyze the envelope of the wave-packet energy as a function of the position along the NPM structure in the direction of transport, as shown in Fig. 4. At non-resonant frequencies, the amplitude of the wave-packed exponentially decays in space. This attenuation is a signature of the anharmonic interaction of the excited phonons with the rest of the phonon population and the boundaries. More precisely, the characteristic length scale of the decay multiplied by the phonon group velocity can be interpreted as the phonon mode relaxation time due to phonon scattering. Interestingly, an exponential relaxation of the wave-packed is not observed either when a resonant frequency is excited in the NPM (see Fig. 4b). This is due to the localization of the mode, which prevents its propagation. These results indicate that the particle-picture of independent phonon evolution based on a velocity and a characteristic anharmonic relaxation time for each mode is insufficient, and highlights the need of integrating the wave and particle pictures of phonons [20, 21, 70] to develop mesoscopic descriptions of heat transport in resonant metamaterials.
@@
We recall that the wave-packet simulations are performed within a thermodynamic environment at 300 K as described by a fully anharmonic interatomic potential accounting for phonon scattering and associated entropy generation [71]. However, the simulated wave-packet is a prescribed mechanical disturbance that selectively excites a frequency range instead of a thermal excitation that perturbs a wide range of phonon modes. This results in an artificial phase-space for phonon interaction where the majority of modes remain close to equilibrium. Moreover, our simulations assume idealized nanostructures with atomically flat surfaces, which promote specular phonon-boundary scattering and enhance the coherent behavior of phonons compared to experimental conditions in samples presenting defects and roughness. Nevertheless, these results explicitly demonstrate that the emergence of phonon mode hybridizations in the band structure, which can be investigated experimentally [56], fundamentally modifies the propagation of thermal phonons even at room temperature. At lower temperatures, the resonant localization effect persists even more strongly as discussed in Appendix VII.4.
@@ V Anderson localization via
By inducing atomic-scale disorder at the length scale of an artifical unit cell instead of the interatomic distance of the crystal’s primitive unit cell, longer wavelength acoustic phonons become prone to localization. For example, the influence of AL has been shown significant in randomally perturbed SLs [15, 16, 18, 19] and NPCs [17]. However, this requires coherent phonon transport to persist over longer distances, and is still limited by the mobility edge frequency. Conversely, since resonance hybridizations influence arbitrary low frequencies, the size of the unit cell in an NPM does not set a lower limit for resonant localization in general. For the small-scale NPM models considered in this work (unit cell size of 3.26 nm), influencing the subwavelength regime enables the localization of low-frequency heat-carrying phonons, which for room-temperature silicon may include wavelengths up to $\sim$10 nm [85]. For longer wavelength phonons that do not accommodate heat capacity, subwavelength resonant localization may take place with arbitrarily large NPM unit cells limited only by the phonon mean free path. Emergence of localized modes in this range can indirectly affect key thermal properties relevant at the nanoscale, such as the non-local length and the heat flux relaxation time which are influenced by a broader distribution of phonons [86]. Another key aspect is that the emergence of phonon interference patterns in AL requires certain stochastic evolution of the non-equilibrium phonon population. In transient conditions at short time scales, such as in the wave-packet simulations, this is thus another limiting factor for energy localization via stochastic interaction with disorder. In contrast, local resonances are established deterministically and instantaneously in NPMs. From this perspective, the presence of local resonances offers a more effective and powerful mechanism than AL for phonon localization.
@@ VI Conclusion
Using frequency-dependent wave-packet excitations introduced into equilibrium MD simulations, where both harmonic and anharmonic effects are active, we have demonstrated explicitly this new phonon localization phenomenon. Parallel simulations of disordered lattices were also conducted to provide direct comparison between the two phonon localization mechanisms. Given that the observed resonant localization phenomenon takes place at discrete narrowband frequency ranges spanning the full spectrum, the impact of the mechanism accumulates leading to fundamental changes to the overall character of the phonon transport. This has implications on the thermal conductivity and, more generally, the broader scope of phonon-related condensed matter phenomena such as phonon-electron couplings and quantum coherence [87, 88].
@@ VII.1.4 Thermal conductivity prediction from equilibrium MD simulations and the Green-Kubo method
We predict and compare the thermal conductivity in the different systems using equilibrium MD simulations followed by application of the G-K method [93]. The systems are initially equilibrated for 1 ns using a time step of size $\Delta{t}=0.5$ fs under the $NPT$ ensemble at 300 K. Then, the heat flux integrated over the volume J is sampled every 4 ps under the $NVE$ ensemble for 6 ns. The resulting flux autocorrelation function $<\textbf{J}(0)\otimes\textbf{J}(t)>$ is used to compute the thermal conductivity

### v2

@@ I Introduction
Distinct from Bragg scattering, a recently discovered mechanism that has been shown to cause transformative changes to the phonon band structure is engineered intrinsic atomic-scale local resonances [39]  [^1] The notion of intrinsic local resonances has been introduced in the context of electromagnetics [99, 100] and acoustics [77, 101, 102]. Unlike thermal phonons, these applications are practically described as linear wave problems where the physical behaviour of interest is dominated by excitations at a single frequency or a narrowband frequency range.. The presence of sub-MFP vibrating substructures generates a multitude of wave-vector-independent resonant phonons (also described as vibrons [41]) that couple and hybridize with the underlying traveling phonons in a host medium [42]. Nanostructures exhibiting this behavior are referred to as nanophononic metamaterials (NPMs); a wide range of NPM configurations have been investigated [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. Unlike the Bragg scatterers in an NPC, the nanoresonators in an NPM are not necessarily periodically spaced, which enhances the robustness of the mechanism. Thermal conductivity reduction by nanoscale local resonances has recently been experimentally demonstrated in suspended silicon membranes with a random array of GaN nanopillars standing on the surface [58]. This nanopillars-on-a-membrane configuration also yielded a decoupling between the thermal and electrical properties, which is highly advantageous for thermolectrics [59, 60].
@@
Here we reveal--for the first time--the appearance of spatial phonon localization along the direction of transport in this class of nanostructures. We provide evidence of this phenomenon by direct explicit observations from room-temperature atomic-scale simulations of silicon membranes with resonating silicon nanopillars standing on the surface. These simulations involve frequency-targeted wave-packet excitations that are applied over a thermalized atomic environment at equilibrium, where the full scope of phonon-phonon scattering events is accounted for  [^2] This contrasts sharply with the localization of linear elastic waves due to local resonances at the macroscopic scale [103].. The form of phonon localization we highlight, which we refer to as resonant localization, occurs in the base membrane along the in-plane direction across multiple unit cells. This is related but distinctly different from the observation of a large concentration of phonon motion along the orthogonal direction in a nanopillar in a single unit cell as described by mode shapes obtained by harmonic eigenvalue analysis--with the latter often also described as a form of localization in the literature [43, 47, 42]. A localized eigen mode is associated with a specific frequency and wavevector pair, whereas the resonant localization that we demonstrate is associated with a single frequency or a narrow band of frequencies and the full scope of all possible wavenumbers along the direction of transport. Resonant localization has profound impact on the thermal behavior and other condensed matter properties in general, and its assessment along the direction of transport enables direct comparison with AL and with coherent phenomena in other types of ordered nanostructures. For example, similar to AL, the localization length or the fraction of localized energy may be characterized. (1) Unlike AL and other coherent phonon mechanisms, this distinct phenomenon manifests across all the phonon spectrum, including low frequencies smaller than the mobility edge of a randomly perturbed lattice (AL) [62, 63, 16] or the Bragg limit corresponding to the periodicity of ordered nanostructure features (SL and NPCs) [42]. (2) Furthermore, unlike AL it is fully deterministically predictable in both space and time. (3) Another unique aspect of the resonant localization mechanism is that it is associated with minimal energy exchange with external sources at the resonant frequency bands.
@@ II Phonon Band Structures
Moving to NPMs, Fig. 1d shows the anharmonic phonon band structure for a resonant unit cell in the form of a membrane with a central squared nanopillar having the same height as the membrane. As seen in previous studies [44, 42], both the LD and MD-SED calculations show in the dispersion diagram the hybridization of numerous phonon modes in this geometry, distinguishable as horizontal lines associated with resonances coupled with the underlying host medium’s wave-vector-dependent propagating modes. The mode shapes of eigenstates populating these horizontal lines (and parts of the associated avoided crossing regions) indicate a concentration of the atomic kinetic energy in the nanoresonator portion of the unit cell, e.g., inside the nanopillar portion of a nanopillared membrane [43, 47, 42, 65]. In contrast to the loss of band gaps in NPCs, band shifts stemming from anharmonicity do not disrupt the presence of the local resonances and their ability to couple with propagating waves, even at room temperature--as shown for the silicon NPM of Fig. 1d. The occurrence of the local resonances may in principle be precisely synthesized to take place across the full phonon spectrum. Indeed, it is also observed in Fig. 1d that the resonance hybridizations take place across the full frequency range displayed, including at low frequencies below the Bragg limit associated with the size of the NPM unit cell. In principle, the number and specific frequencies of these resonances are a characteristic of the nanopillar geometry and choice of material, and are therefore controllable by design. The number of local resonances is three times the number of atoms in a unit nanoresonator, and the bandwidth of each resonant mode is dependent on its dissipation properties [47]. Anharmonic unit-cell analysis may also be done by lattice dynamics-based calculations; for example the quasiharmonic Green-Kubo method [21] was used to predict a frequency-dependent diffusivity quantity corresponding to the anharmonic phonon band-structure of an NPM unit cell [50]
@@ III Deterministic localization via local resonances
Figure 4: Wave-packet envelope evolution at room temperature. Maximum phonon kinetic energy measured around the intervention region resulting from wave-packet excitation at the resonant frequency (b) or non-resonant frequencies (a,c) in NPM. The kinetic energy is represented in logarithmic scale, and the dashed lines indicate exponential fits to the envelope relaxation. As shown in (b), this fit is unsuccessful under the resonant conditions due to the lack of wave front propagation, thus indicating unusual phonon propagation and relaxation beyond the particle-like scattering picture provided by the relaxation time approximation.
@@
To further explore the resonant localization phenomenon, it is illustrative to analyze the envelope of the wave-packet energy as a function of the position along the NPM structure in the direction of transport, as shown in Fig. 4. At non-resonant frequencies, the amplitude of the wave packet exponentially decays in space. This attenuation is a signature of the anharmonic interaction of the excited phonons with the rest of the phonon population and the boundaries. Nonlinear transfer of wave-packet energy to other phonon modes dissipates the energy accommodated in the wave front into the thermal background, thus leading to diffusive evolution at long times. Precisely, the characteristic length scale of the decay multiplied by the phonon group velocity can be interpreted as the phonon mode relaxation time due to phonon scattering. This interpretation is confirmed by wave-packet simulations at 0 K, where the lack of anharmonic effects prevents the decay of the wave-packet amplitude. A key observation is that an exponential relaxation of the wave-packed is similarly not observed when a resonant frequency is excited in the NPM at 300 K. In this case, the excited phonon modes are fully localized, and the only available mechanism to relax the perturbation is the nonlinear transfer of energy to modes with non-resonant frequencies that can propagate. In Figure Fig. 4b, this dissipative relaxation process translates to both a lack of wave fronts and a non-exponential decay of the wave-packet envelope. These effects are thus unambiguous signatures of deterministic localization that can be qualitatively distinguished from velocity reduction effects due to band flattening. In turn, these results involving resonances indicate that the particle-picture of independent phonon evolution based on a velocity and a characteristic anharmonic relaxation time for each mode is insufficient, and highlights the need of integrating the wave and particle pictures of phonons [20, 21, 71] to develop mesoscopic descriptions of heat transport in resonant metamaterials.
@@
We recall that the wave-packet simulations are performed within a thermodynamic environment at 300 K as described by a fully anharmonic interatomic potential accounting for phonon scattering and associated entropy generation [72]. However, the simulated wave packet is a prescribed mechanical disturbance that selectively excites a frequency range instead of a thermal excitation that perturbs a wide range of phonon modes. This results in an artificial phase-space for phonon interaction where the majority of modes remain close to equilibrium. Moreover, our simulations assume idealized nanostructures with atomically flat surfaces, which promote specular phonon-boundary scattering and enhance the coherent behavior of phonons compared to experimental conditions in samples presenting defects and roughness. Nevertheless, these results explicitly demonstrate that the emergence of phonon mode hybridizations in the band structure fundamentally modifies the propagation of thermal phonons even at room temperature. These results will further guide experimental research, noting that there are cases where the impact of the resonances has been observed, e.g., on the thermal conductivity [58] and on the low-frequency band structure [56], but also cases where such characterization was not possible [73, 74, 75] [^3] Among the experimental challenges for NPM characterization are: (1) the presence of fabrication defects or imperfections, which enhance the role of anharmonicity at room temperature relative to atomically-perfect crystals, (2) fabrication of prototypes with a very low pillar-to-membrane volume ratio, which weakens the impact of the local resonances, and (3) examination of configurations where the unit cell is overly large and comparable to the mean free path, which weakens the degree of presence of coherent phonons. At lower temperatures, the resonant localization effect persists even more strongly as discussed in Appendix VII.4.
@@ V Anderson localization via
By inducing atomic-scale disorder at the length scale of an artifical unit cell instead of the interatomic distance of the crystal’s primitive unit cell, longer wavelength acoustic phonons become prone to localization. For example, the influence of AL has been shown significant in randomally perturbed SLs [15, 16, 18, 19] and NPCs [17]. However, this requires coherent phonon transport to persist over longer distances, and is still limited by the mobility edge frequency. Conversely, since resonance hybridizations influence arbitrary low frequencies, the size of the unit cell in an NPM does not set a lower limit for resonant localization in general. For the small-scale NPM models considered in this work (unit cell size of 3.26 nm), influencing the subwavelength regime enables the localization of low-frequency heat-carrying phonons, which for room-temperature silicon may include wavelengths up to $\sim$10 nm [90]. For longer wavelength phonons that do not accommodate heat capacity, subwavelength resonant localization may take place with arbitrarily large NPM unit cells limited only by the phonon mean free path. Emergence of localized modes in the low-frequency range can indirectly affect key thermal properties relevant at the nanoscale, such as the non-local length and the heat flux relaxation time--which are influenced by a broader distribution of phonons compared to the range that directly influences the thermal conductivity [91]. Another key aspect is that the emergence of phonon interference patterns in AL requires certain stochastic evolution of the non-equilibrium phonon population. In transient conditions at short time scales, such as in the wave-packet simulations, this is thus another limiting factor for energy localization via stochastic interaction with disorder. In contrast, local resonances are established deterministically and instantaneously in NPMs. From this perspective, the presence of local resonances offers a more effective and powerful mechanism than AL for phonon localization.
@@ VI Conclusion
Using frequency-dependent wave-packet excitations introduced into equilibrium MD simulations, where both harmonic and anharmonic effects are active, we have demonstrated explicitly this new phonon localization phenomenon. Parallel simulations of disordered lattices were also conducted to provide direct comparison between the two phonon localization mechanisms. Given that the observed resonant localization phenomenon takes place at discrete narrowband frequency ranges spanning the full spectrum, and highly tunable by design, the impact of the mechanism accumulates leading to fundamental changes to the overall character of the phonon transport. This has implications on the thermal conductivity and, more generally, the broader scope of phonon-related condensed matter phenomena such as phonon-electron couplings and quantum coherence [92, 93].
@@ VII.1.4 Thermal conductivity prediction from equilibrium MD simulations and the Green-Kubo method
We predict and compare the thermal conductivity in the different systems using equilibrium MD simulations followed by application of the G-K method [98]. The systems are initially equilibrated for 1 ns using a time step of size $\Delta{t}=0.5$ fs under the $NPT$ ensemble at 300 K. Then, the heat flux integrated over the volume J is sampled every 4 fs under the $NVE$ ensemble for 6 ns. The resulting flux autocorrelation function $<\textbf{J}(0)\otimes\textbf{J}(t)>$ is used to compute the thermal conductivity
@@ VII.2 Fraction of Localized Energy
As mentioned earlier, an attractive aspect of resonant localization is the ability to access exceedingly low frequencies, e.g., below the mobility edge that limits AL. This is beneficial for multiple reasons. First, it allows for a stronger reduction in the thermal conductivity (without the need to increase the unit-cell size) because low-frequency phonons contribute strongly to the thermal conductivity. Second, other average quantities beyond the thermal conductivity, such as the non-local length characterizing the heat flux correlations [91], are determined by a broader spectrum than the thermal conductivity. Third, other phenomenon, such as phonon-electron coupling for example, are also generally influenced by a broader portion of the spectrum including low frequencies.
Another favorable feature of resonant localization is its high degree of tunability. It is possible to determine a priori the precise spatial distribution of the localization by controlling the locations where the nanoresonators are placed within the crystal. In addition, the target frequencies for localization exactly correspond to the resonances of the nanopillar, which can be explicitly computed.